OpenStudy (anonymous):

Find the values of x and y that maximize or minimize the objective function for each of these graphs. a)Minimum for P = 3x + y B) maximum for p= x+ y

OpenStudy (anonymous):

a

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@freckles please help

OpenStudy (freckles):

this is linear programming?

OpenStudy (freckles):

i think it is take your corner points and plug them in

OpenStudy (anonymous):

i dont have a calculator

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

you don't need a calculator

OpenStudy (freckles):

they even tell you the points for the corners in your pictures

OpenStudy (freckles):

for the first picture I see (0,5) and (3,0) and (2,1) plug those in to your first equation

OpenStudy (freckles):

for example P(x,y)=3x+y P(0,5)=3*0+5=? and then so on you will plug all three of those points in and see which gives you the smallest number

OpenStudy (freckles):

since you are looking for min value in the first problem

OpenStudy (anonymous):

i just really dont understand this

OpenStudy (freckles):

can you not do the operations I suggested above?

OpenStudy (freckles):

P(0,5)=3*0+5=?

OpenStudy (anonymous):

=5

OpenStudy (anonymous):

?

OpenStudy (freckles):

then plug in the other points

OpenStudy (freckles):

P(2,1)=3*2+1=?

OpenStudy (anonymous):

7

OpenStudy (freckles):

and plug in the last point that was given

OpenStudy (freckles):

then compare the outputs

OpenStudy (anonymous):

p(0,3)=3*0+1?

OpenStudy (freckles):

(3,0) isn't the same as (0,3)

OpenStudy (anonymous):

0*3?

OpenStudy (freckles):

plug in (3,0) not (0,3)

OpenStudy (anonymous):

i really dont get it

OpenStudy (anonymous):

i tried both ways

OpenStudy (freckles):

no you haven't plugged in (3,0) which is point they want you to plug in they never said anything about (0,3)

OpenStudy (anonymous):

p(3,0)=3*0+1?

OpenStudy (freckles):

P(x,y)=3x+y (3,0) is not the same as (0,3) (3,0) means x is 3 and y is 0 (0,3) means x is 0 and y is 3 <---we aren't suppose to plug this in

OpenStudy (freckles):

you said you wanted to plug in (3,0) but you plug in (0,1)

OpenStudy (anonymous):

3+1?

OpenStudy (freckles):